OpenStudy (anonymous):

can anyone solve this. 3x=x+20? answer please..

OpenStudy (anonymous):

x =10

OpenStudy (anonymous):

x=-2, unless its not a 0 at the end.

OpenStudy (anonymous):

show me the solution

OpenStudy (anonymous):

3x=x+20 subtract x on both sides 3x-x = x+20-x 2x = 20 divide by 2 on both sides x =10

OpenStudy (anonymous):

checking? i dunno how to check the answer

OpenStudy (anonymous):

plug in 10 where ever you see x and see if LHS = RHS